# HI-TransPA: Hearing Impairments Translation Personal Assistant

### v1

@@ Abstract
To provide a unified and flexible solution for daily communication among hearing‑impaired individuals, we introduce the Omni‑Model paradigm into assistive technology and present HI‑TransPA, an instruction‑driven audio‑visual personal assistant. The model fuses indistinct speech with high‑frame‑rate lip dynamics, enabling both translation and dialogue within a single multimodal framework. To tackle the challenges of noisy and heterogeneous raw data and the limited adaptability of existing Omni‑Models to hearing‑impaired speech, we construct a comprehensive preprocessing and curation pipeline that detects facial landmarks, isolates and stabilizes the lip region, and quantitatively assesses multimodal sample quality. These quality scores guide a curriculum learning strategy that first trains on clean, high‑confidence samples and progressively incorporates harder cases to strengthen model robustness. We further adopt a SigLIP encoder combined with a Unified 3D‑Resampler to efficiently encode high‑frame‑rate lip motion. Experiments on our purpose‑built HI‑Dialogue dataset show that HI‑TransPA achieves state‑of‑the‑art performance in both literal accuracy and semantic fidelity. This work establishes a foundation for applying Omni‑Models to assistive communication technology, providing an end‑to‑end modeling framework and essential processing tools for future research.
@@ 1 Introduction
Despite the remarkable progress of artificial intelligence, accessibility and real-world applicability remain limited for certain groups. Individuals with disabilities constitute a substantial share of the global population and still encounter barriers that prevent them from fully benefiting from AI technologies. Encouragingly, early research efforts have begun to bridge this gap by demonstrating how AI can promote greater inclusivity. For instance, vision-based systems such as “electronic guide dogs” [InternDog] have been developed to assist visually impaired users for their mobility and independence. These pioneering works underscore the tremendous potential of socially oriented, human-centric AI in promoting accessibility.
Most existing assistive technologies focus on converting speech from hearing individuals into text, enabling deaf or hard-of-hearing users to access spoken information. However, these systems offer little support when hearing-impaired users attempt to express themselves. Conventional speech recognition models are typically trained on standard speech data and therefore struggle to interpret atypical or partially articulated utterances. This limitation hinders the ability of hearing-impaired individuals to engage in spoken interaction and underscores the need for adaptive multimodal models that can robustly understand and process diverse and non-standard speech characteristics.
To this end, we introduce the Omni-Model paradigm for assistive technologies targeting hearing-impaired users and present HI-TransPA, an instruction-driven audio-visual personal assistant. To the best of our knowledge, this is the first unified framework that integrates translation and dialogue functionalities within a single model explicitly designed for assisting the hearing-impaired. HI-TransPA simultaneously processes both raw audio (e.g., indistinct or impaired speech) and high-frame-rate visual cues derived from lip dynamics, employing instruction tuning to perform diverse communication tasks with enhanced flexibility. Beyond functioning as a translator, it acts as an intelligent conversational partner that understands context, infers user intent, and generates appropriate responses. Furthermore, to catalyze progress in accessible multimodal AI, we open-source an automated data-processing framework for extracting and aligning lip-sequence features from real-world recordings.
@@
    Unified audio‑visual Omni‑Model. We introduce HI‑TransPA, the first instruction‑driven audio‑visual Omni‑Model for the hearing‑impaired, unifying translation and dialogue within a single multimodal framework.
@@
    Cross‑modal comprehension. HI‑TransPA jointly models auditory signals and high‑frame‑rate lip dynamics via instruction tuning, achieving robust cross‑modal alignment and context‑aware understanding in complex real‑world scenarios.
@@
    Quality‑guided curriculum learning. A data‑driven curriculum strategy leverages multimodal quality scoring to schedule training from easy to hard samples, substantially improving generalization to diverse hearing‑impaired speech.
@@ 2.1 ASR and Multimodal Large Language Models
The field of audio-centric AI has rapidly evolved from narrow, task-specific systems to comprehensive multimodal understanding. Early progress in Automatic Speech Recognition (ASR) was driven by large-scale weakly supervised training (e.g., Whisper [whisper]), data-efficient modeling (e.g., Canary [canary]), and architectures optimized for fast inference (e.g., Paraformer [paraformer]). A major paradigm shift subsequently occurred with the integration of Large Language Models (LLMs), as demonstrated by FireRedASR [fireredasr], which enhances ASR performance through greater contextual comprehension and reasoning.
@@
Building upon these advances, Omni-Models have emerged to unify text, audio, and visual modalities within a single foundation model. Both proprietary and open-source systems, including GPT-4o [gpt], Qwen2.5-Omni [qwen2omni], and MiniCPM-o 2.6, demonstrate impressive multimodal reasoning capabilities. However, their applications in assistive technology for individuals with hearing loss remain limited. Progress toward inclusive communication tools is further constrained by the scarcity of specialized multimodal datasets, particularly those containing indistinct speech aligned with lip dynamics.
@@ 3.1 Data Preprocessing and Curation
In the first stage, for each video frame $I_{t}$, a pre-trained landmark detector $F_{lm}$ extracts $N=468$ facial keypoints, from which we retain only the $M$ points corresponding to the lip region:
@@
In the second stage, we align and stabilize the cropped lip video $V_{\text{lips}}$. For frames with valid landmarks $P_{t}$, we compute bounding boxes $B_{t}$ and define a uniform crop size $S$ as:
@@
where $\gamma=1.2$ is an expansion factor and $\lceil\cdot\rceil_{2}$ rounds up to the nearest even number. Cropping is centered at each frame’s landmark centroid $C_{t}$, linearly interpolated when landmarks are missing. The resulting stabilized video $V_{\text{lips}}$ normalizes head motion and emphasizes lip dynamics relevant to speech articulation.
@@
Even after region extraction, hearing-impaired speech data may still suffer from articulation ambiguity or acoustic distortion. To ensure data reliability, we introduce a rejection sampling framework that scores each audio-visual pair and partitions the dataset into accepted and rejected subsets, as shown in Fig. 4.
@@
$S_{\text{audio}}(x)=0.5\cdot S_{\text{ASR}}(x)+0.5\cdot\text{Norm}(SNR(x)).$ (5)
@@ 3.3 Multi-Stage Alignment and Fine-Tuning
Finally, we conduct end-to-end fine-tuning on our HI-Dialogue dataset using the curriculum learning strategy (Section 3.4). This step enables the model to adapt to natural conversational scenarios of hearing-impaired users.
@@ 3.4 Training Objective
$\mathcal{L}_{\text{Stage 1}}=\mathbb{E}_{x\in\mathcal{D}_{\text{accept}}}[L_{\text{CE}}(f(x),y)].$ (10)
This stage establishes stable multimodal alignment between lip motion, audio, and textual content.
@@
$\mathcal{L}_{\text{Stage 2}}=\mathbb{E}_{x^{\prime}\in\mathcal{D}_{\text{reject}}}[L_{\text{CE}}(f(x^{\prime}),y^{\prime})].$ (11)
This step implicitly up-weights challenging examples, encouraging the model to learn robust representations that generalize to noisy, real-world conditions.
@@
For evaluation, we report the Character Error Rate (CER) to measure literal transcription accuracy. Given a ground-truth sequence $y$ and model output $\hat{y}$, CER is defined as:
@@ 4 Experiments
Models Parameters Audio Video EmbSim CER(%) CS
@@
Whisper-large-V3 1.6B $\checkmark$ 0.79 0.32 0.74
SenseVoice-small 0.3B $\checkmark$ 0.71 0.35 0.68
Paraformer-large 0.2B $\checkmark$ 0.70 0.38 0.66
FireRedASR 8.3B $\checkmark$ 0.77 0.38 0.70
@@
Qwen2-Audio 7B $\checkmark$ 0.74 0.44 0.65
MiDashengLM 7B $\checkmark$ 0.67 0.53 0.57
InternLM-XComposer2.5-OmniLive 7B $\checkmark$ 0.67 0.54 0.57
Step-Audio 2 mini 7B $\checkmark$ 0.79 0.34 0.73
@@
Qwen2.5-Omni (3B) 3B $\checkmark$ $\checkmark$ 0.73 0.44 0.65
Qwen2.5-Omni (7B) 7B $\checkmark$ $\checkmark$ 0.75 0.42 0.67
HI-TransPA 3B $\checkmark$ $\checkmark$ 0.77 0.37 0.70
HI-TransPA (Curriculum Learning) 3B $\checkmark$ $\checkmark$ 0.84 0.27 0.79
@@ 4.1 Experimental Setup
Dataset. We collected and curated a dedicated dataset, HI-Dialogue, to support the training and evaluation of audio-visual dialogue models for hearing-impaired individuals. Six volunteers with varying levels of hearing loss recorded audio-visual materials covering daily conversations, instructional texts, and emergency situations. After manual screening to remove samples with occluded lip regions or mismatched transcripts, we obtained 9,763 high-quality samples.
@@ 4.3 Baselines
ASR Models. We select four widely used automatic speech recognition (ASR) systems representing strong audio-only foundations with different design philosophies. Whisper-large-v3 [whisper] is trained on 680k hours of weakly supervised web data and serves as a robust multilingual reference under zero-shot conditions. SenseVoice-small [sensevoice] is part of the FunAudioLLM framework, supporting fast multilingual ASR and audio-event recognition with low latency. Paraformer-large [paraformer] employs a non-autoregressive parallel Transformer to accelerate inference while maintaining competitive accuracy. FireRedASR [fireredasr] represents an industrial-grade, Mandarin-optimized encoder–decoder architecture integrating large language modeling for superior character error rate (CER) performance. These models quantify the upper bound of purely acoustic transcription capability and provide a foundation for measuring the gains from incorporating visual inputs.
@@
Omni-Models. Finally, we compare with open-source multimodal Omni-Models that jointly process audio, visual, and textual information. Qwen2.5-Omni (3B and 7B) [qwen2omni] integrates synchronized audio–visual encoders and a shared language backbone to handle streaming multimodal tasks, representing the state of general-purpose cross-modal LLMs. Our HI-TransPA and its variant trained with curriculum learning follow the Omni-Model paradigm but are specifically optimized for accessibility. By aligning high-frame-rate lip features and noisy audio inputs through instruction tuning, HI-TransPA demonstrates notable improvements in both embedding similarity and character error rate, outperforming generic Omni baselines by a clear margin.
@@ 4.4 Main Results and Analysis
Results are presented in Table 1.
@@ 4.5 Interpretation of the Comprehensive Score
Understanding what CS reveals about model comprehension performance. The comprehensive score (Eq. 14) jointly captures literal correctness $(1-\text{CER})$ and semantic coherence (EmbSim), enabling us to compare models beyond raw recognition accuracy.
From Table 1, three core observations emerge:
@@
    Generic Omni-models: Qwen2.5-Omni (3B and 7B) achieve balanced yet limited improvement. They benefit from visual cues but fail to capture fine-grained temporal lip dynamics, keeping both CER and EmbSim suboptimal.
@@
    HI-TransPA: Our model enhances both dimensions simultaneously—reducing CER from 42% to 27% while raising EmbSim from 0.75 to 0.84—leading to CS = 0.79. This implies that HI-TransPA not only produces more accurate transcripts but also better preserves contextual meaning through specialized lip-motion modeling and curriculum-based training.
@@ 4.6 Ablation Studies
4.7 Qualitative Analysis
To provide a more intuitive understanding of our model’s comprehension capabilities and how they directly impact the quality of downstream dialogue, we present two representative cases in Figure 6.
In the “Good case,” our HI-TransPA model successfully processes a complex sentence containing multiple entities. The output in “Translate mode” is a perfect transcription of the original content, demonstrating its robust multimodal understanding. Based on this precise comprehension, the “Chat mode” is able to generate a highly relevant, insightful, and empathetic response, truly functioning as an intelligent assistant.
In contrast, the “Bad case” vividly illustrates the consequences of comprehension failure (this result is from a poorly performing baseline). The model completely misinterprets the user’s intent in “Translate mode,” producing a sentence that is entirely unrelated to the ground truth. This upstream failure in comprehension leads to a catastrophic downstream result in “Chat mode,” where the model generates a nonsensical dialogue based on the erroneous text.
This failure case powerfully validates our Comprehension First principle, as outlined in Section 4.2. It clearly shows that when a model fails at the critical comprehension stage, any subsequent dialogue generation becomes meaningless and potentially misleading, regardless of the power of its internal language model. This highlights how HI-TransPA, through its superior comprehension abilities as quantified in Table 1, lays a solid foundation for providing truly meaningful and helpful conversational support.

### v2

@@ Abstract
Hearing-impaired individuals often face significant barriers in daily communication due to the inherent challenges of producing clear speech. To address this, we introduce the Omni-Model paradigm into assistive technology and present HI-TransPA, an instruction-driven audio-visual personal assistant. The model fuses indistinct speech with lip dynamics, enabling both translation and dialogue within a single multimodal framework. To address the distinctive pronunciation patterns of hearing-impaired speech and the limited adaptability of existing models, we develop a multimodal preprocessing and curation pipeline that detects facial landmarks, stabilizes the lip region, and quantitatively evaluates sample quality. These quality scores guide a curriculum learning strategy that first trains on clean, high-confidence samples and progressively incorporates harder cases to strengthen model robustness. Architecturally, we employs a novel unified 3D-Resampler to efficiently encode the lip dynamics, which is critical for accurate interpretation. Experiments on purpose-built HI-Dialogue dataset show that HI-TransPA achieves state-of-the-art performance in both literal accuracy and semantic fidelity. Our work establishes a foundation for applying Omni-Models to assistive communication technology, providing an end-to-end modeling framework and essential processing tools for future research.
@@ 1 Introduction
Despite the remarkable progress of artificial intelligence, accessibility and real-world applicability remain limited for certain groups. Individuals with disabilities constitute a substantial share of the global population and still encounter barriers that prevent them from fully benefiting from AI technologies. Encouragingly, early research efforts have begun to bridge this gap by demonstrating how AI can promote greater inclusivity. For instance, vision-based systems such as electronic guide dogs [InternDog] and OpenAIglasses for Navigation [AI-FanGe2025OpenAIglasses] have been developed to assist visually impaired users in enhancing their mobility and independence. These pioneering works underscore the potential of socially oriented, human-centric AI in promoting accessibility.
Most existing assistive technologies focus on converting speech from hearing individuals into text, enabling deaf or hard-of-hearing users to access spoken information. However, these systems offer little support when hearing-impaired users attempt to express themselves. Conventional speech recognition models [sensevoice, paraformer] are typically trained on standard speech data and therefore struggle to interpret atypical or partially articulated utterances. This limitation hinders the ability of hearing-impaired individuals to engage in spoken interaction and underscores the need for adaptive multimodal models that can understand and process diverse and non-standard speech characteristics.
@@
To this end, we introduce the Omni-Model paradigm for assistive technologies targeting hearing-impaired users and present HI-TransPA, an instruction-driven audio-visual personal assistant. To the best of our knowledge, this is the first unified framework that integrates translation and chat functionalities within a single model explicitly designed for assisting the hearing-impaired. To enable HI-TransPA to learn effectively from such inherently challenging data, we first design a comprehensive data processing pipeline that purifies input signals and then evaluates and partitions the data based on quality. Building upon this partitioning, we then propose a novel quality-aware curriculum learning strategy that leverages the structured data for robust training. Through these methods, HI-TransPA simultaneously processes both raw audio and high-frame-rate visual cues from lip dynamics, and acts as an intelligent conversational partner that understands context and infers user intent.
@@
    Unified audio-visual framework. We introduce HI-TransPA, an instruction-driven multimodal framework for assisting hearing-impaired individuals. By jointly modeling auditory signals and high-frame-rate lip dynamics, HI-TransPA performs accurate translation and natural dialogue within a single end-to-end architecture.
@@
    Comprehensive multimodal data pipeline. We construct a full pipeline covering preprocessing and quality-aware curation. It first extracts and aligns the lip region through a two-stage method, then applies rejection sampling based on multimodal metrics to automatically split data into “easy” and “hard” subsets.
@@
    Quality-aware curriculum learning. We propose a data-driven curriculum strategy that ranks multimodal samples by quality and schedules training from easy to hard, leading to stronger generalization across diverse hearing-impaired speech patterns.
@@ 2.1 ASR and Multimodal Large Language Models
The field of audio-centric AI has rapidly evolved from narrow, task-specific systems to comprehensive multimodal understanding. Early progress in Automatic Speech Recognition (ASR) was driven by large-scale weakly supervised training (e.g., Whisper [whisper]), data-efficient modeling (e.g., Canary [canary]), and architectures optimized for fast inference (e.g., Paraformer [paraformer]). A major paradigm shift subsequently occurred with the integration of Large Language Models (LLMs), as demonstrated by FireRedASR-AED [fireredasr], which enhances ASR performance through greater contextual comprehension and reasoning.
@@
Building upon these advances, Omni-Models have emerged to unify text, audio, and visual modalities within a single foundation model. Both proprietary and open-source systems, including GPT-4o [gpt], and Qwen2.5-Omni [qwen2omni], demonstrate impressive multimodal reasoning capabilities. However, their applications in assistive technology for individuals with hearing loss remain limited. Progress toward inclusive communication tools is further constrained by the scarcity of multimodal datasets, particularly those that containing indistinct speech aligned with lip dynamics.
@@ 3.1 Data Preprocessing and Curation
In the first stage, for each video frame $I_{t}$, a pre-trained landmark detector $F_{lm}$ extracts $N=468$ facial keypoints with coordinates $(x_{i},y_{i})$, from which we retain only the $M$ points corresponding to the lip region:
@@
In the second stage, we align and stabilize the lip video $V_{\text{lips}}$. For frames with valid landmarks $P_{t}$, we compute bounding boxes $B_{t}$ and define a uniform crop size $S$ as:
@@
where $\gamma=1.2$ is chosen as expansion factor and $\lceil\cdot\rceil_{2}$ rounds up to the nearest even number. Cropping is centered at each frame’s landmark centroid $C_{t}$, linearly interpolated when landmarks are missing. The resulting stabilized video $V_{\text{lips}}$ normalizes head motion and emphasizes lip dynamics relevant to speech articulation.
@@
Even after region extraction, hearing-impaired speech data can exhibit articulation ambiguity or acoustic distortion. To ensure data reliability, we introduce a rejection sampling framework that scores each audio-visual pair and partitions the dataset into accepted and rejected subsets, as shown in Fig. 4.
@@
$S_{\text{audio}}(x)=0.5\cdot S_{\text{ASR}}(x)+0.5\cdot\text{Norm}(SNR(x)),$ (5)
where $\text{Norm}(\cdot)$ denotes min-max normalization.
@@ 3.3 Multi-Stage Alignment and Fine-Tuning
We perform end-to-end instruction fine-tuning to jointly optimize the model’s translation and dialogue abilities. A mixed-instruction dataset is constructed with two modes: /translate, using paired audio-visual samples and their reference texts; and /chat, which shares the same inputs but pairs them with language-model–generated conversational replies. During training, both modes are combined under the curriculum learning strategy described in Section 3.4, where sample difficulty is determined by translation quality. The unified fine-tuning stage enables HI-TransPA to consolidate comprehension and transcription skills while adapting to real conversational scenarios.
@@ 3.4 Training Objective
$\mathcal{L}_{\text{Stage 1}}=\mathbb{E}_{x\in\mathcal{D}_{\text{accept}}}[L_{\text{CE}}(f(x),y)],$ (10)
which aim to establish stable multimodal alignment between lip motion, audio, and textual content.
@@
$\mathcal{L}_{\text{Stage 2}}=\mathbb{E}_{x^{\prime}\in\mathcal{D}_{\text{reject}}}[L_{\text{CE}}(f(x^{\prime}),y^{\prime})],$ (11)
which is designed to implicitly up-weights challenging examples, encouraging the model to learn robust representations that generalize to noisy, real-world conditions.
@@ 3.5 Metrics Definition
For evaluation, we report the CER to measure literal transcription accuracy. Given a ground-truth sequence $y$ and model output $\hat{y}$, CER is defined as:
@@ 4 Experiments
Models Params Audio Video EmbSim CER CS
@@
Whisper-large-V3 1.6B 0.79 0.32 0.74
SenseVoice-small 0.3B 0.71 0.35 0.68
Paraformer-large 0.2B 0.70 0.38 0.66
FireRedASR-AED 1.1B 0.77 0.38 0.70
@@
Qwen2-Audio 7B 0.74 0.44 0.65
MiDashengLM 7B 0.67 0.53 0.57
InternLM-XComposer2.5-OmniLive 7B 0.67 0.54 0.57
Step-Audio 2 mini 7B 0.79 0.34 0.73
@@
Qwen2.5-Omni (3B) 3B 0.73 0.44 0.65
Qwen2.5-Omni (7B) 7B 0.75 0.42 0.67
HI-TransPA 3B 0.77 0.37 0.70
HI-TransPA (Curriculum Learning) 3B 0.84 0.27 0.79
@@ 4.1 Experimental Setup
Dataset. We collected and curated a dedicated dataset, HI-Dialogue, to support the training and evaluation of audio-visual dialogue models for hearing-impaired individuals. Six volunteers with varying levels of hearing loss recorded audio-visual materials covering daily conversations, instructional texts, and emergency situations. After manual screening to remove samples with occluded lip regions or mismatched transcripts, we obtained 9,673 high-quality samples.
@@ 4.3 Baselines
ASR Models. We select four widely used automatic speech recognition (ASR) systems representing strong audio-only foundations with different design philosophies. Whisper-large-v3 [whisper] is trained on 680k hours of weakly supervised web data and serves as a robust multilingual reference under zero-shot conditions. SenseVoice-small [sensevoice] is part of the FunAudioLLM framework, supporting fast multilingual ASR and audio-event recognition with low latency. Paraformer-large [paraformer] employs a non-autoregressive parallel Transformer to accelerate inference while maintaining competitive accuracy. FireRedASR-AED [fireredasr] represents an industrial-grade, Mandarin-optimized encoder–decoder architecture integrating large language modeling for superior character error rate (CER) performance. These models quantify the upper bound of purely acoustic transcription capability and provide a foundation for measuring the gains from incorporating visual inputs.
@@
Omni-Models. Finally, we compare with open-source multimodal Omni-Models that jointly process audio, visual, and textual information. Qwen2.5-Omni (3B and 7B) [qwen2omni] integrates synchronized audio–visual encoders and a shared language backbone to handle streaming multimodal tasks, representing the state of general-purpose cross-modal LLMs. Our HI-TransPA and its variant trained with curriculum learning follow the Omni-Model paradigm but are specifically optimized for accessibility. By aligning high-frame-rate lip features and audio inputs through instruction tuning, HI-TransPA demonstrates improvements in both embedding similarity and character error rate, outperforming generic Omni baselines by a clear margin.
@@ 4.4 Main Results and Analysis
The detailed comparison results are presented in Table 1.
@@ 4.5 Interpretation of the Comprehensive Score
The comprehensive score (Eq. 14) combines literal correctness $(1-\text{CER})$ and semantic coherence (EmbSim), enabling us to compare models beyond recognition accuracy. From Table 1, three core observations emerge:
@@
    Generic Omni-models: Qwen2.5-Omni (3B and 7B) achieve modest improvement. They benefit from visual cues but fail to capture fine-grained temporal lip dynamics, keeping both CER and EmbSim suboptimal.
@@
    HI-TransPA: Our model enhances both dimensions simultaneously, reducing the CER from 42% to 27% and increasing the EmbSim from 0.75 to 0.84, which leads to a CS of 0.79. This indicates that HI-TransPA not only produces more accurate transcripts but also better preserves contextual meaning through specialized lip-motion modeling and curriculum-based training.
@@ 4.6 Ablation Studies
4.7 Qualitative Analysis
To provide a more intuitive demonstration of our model’s capabilities and to analyze the critical role of comprehension accuracy in generating meaningful dialogue, we present two representative cases in Figure 6. These cases showcase two modes of operation for HI-TransPA: a “Translate mode,” which aims for a literal transcription that reveals the model’s true understanding of the input, and a “Chat mode,” which aims to generate a natural and helpful conversational response based on that understanding.
@@ 4.7 Qualitative Analysis
In the “Good case,” our HI-TransPA model successfully processes a complex sentence containing multiple entities. The output in “Translate mode” is a near-perfect transcription, demonstrating its robust multimodal comprehension. Based on this precise understanding, the “Chat mode” is able to go beyond simple repetition, infer the broader context of an “international grain shipment,” and provide an empathetic and insightful response, truly functioning as an intelligent assistant.
In contrast, the “Bad case” vividly illustrates what occurs when comprehension fails (this result is from a poorly performing baseline). The “Translate mode” output reveals a complete misinterpretation of the user’s intent, producing a sentence semantically unrelated to the original input about transferring money at a bank. This fundamental failure in comprehension directly explains why the “Chat mode” response is nonsensical in context. Although the chat response is grammatically coherent, it is logically derived from an incorrect premise, rendering it entirely useless and potentially confusing for the user.
The comparison strongly validates our Comprehension First principle (Section 4.2), showing that accurate comprehension, reflected in literal translation, forms the basis of all higher-level dialogue abilities. As confirmed in Table 1, HI-TransPA’s superior understanding establishes a solid foundation for meaningful and supportive interaction.